I am lost on how to work this out!! Can you please explain step by step?
Dave Sharp has $10 to buy stamps. He wants to buy twice as many .22 cent stamps as .14 cent stamps. What is the greatest number of .22 cent stamps he can buy?

Answer by ilana(305)

I would just do guess and check. Every time he buys another 2 $.22 stamps, he buys 1 more $.14 stamp. So he spends .22+.22+.14=$.58 at a time. This might need to be changed when we get closer to the answer, but for now this works.
So let's start with 10 of these bunches- that's $5.80. Slightly more than half. Try 18- that's $10.44- a bit too much! Try 17- 9.86- so it's around 17 bunches. Now, we had 2 $.22 stamps in each, so that's about 34. So if it was 34 $.22, it's 17 $.14 stamps. Does that work? (34*.22)+(17*.14)=7.48+2.38=9.86- so that part worked. If he buys one more of anything, he goes over. So the answer must be 34.